1b) Two cards are drawn simultaneously or successively without replacement from a

well-shuffled deck of 52 cards. Find the probability distribution of number of aces

).x( Give a graphical representation of probability distribution


1
Expert's answer
2021-05-07T10:28:42-0400

Solution:

Let X denote the number of aces in a sample of 2 cards drawn from a well-shuffled pack of 52 playing cards.

So, X can take the values 0,1 and 2.

Now,

"P(X=0)
\\
=P( no\ ace )

=\frac{48}{52} \times \frac{47}{51}
\\
=\frac{2256}{2652}

=\frac{188}{221}"

"P(X=1)
\\
=P(1\ ace ) =\frac{4}{52} \times \frac{48}{51}+\frac{48}{52} \times \frac{4}{51}
\\
=\frac{384}{2652}
=\frac{32}{221}"

"P(X=2)\\
=P(2\ aces )

=\frac{4}{52} \times \frac{3}{51}\\
=\frac{12}{2652}
=\frac{1}{221}\\"

"\text { Thus, the probability distribution of } X \text { is given by }\\"